COMPLETE EMBEDDED MINIMAL

SURFACES OF FINITE TOTAL CURVATURE

David Hoffman Hermann Karcher

Abstract. We survey what is known about minimal surfaces in R^{3} that are complete, embedded, and have finite total curvature. The only classically known examples of such surfaces were the plane and the catenoid. The discovery by Costa, early in the last decade, of a new example that proved to be embedded sparked a great deal of research in this area. Many new examples have been found, even families of them, as will be described below. The central question has been transformed from whether or not there are any examples except surfaces of rotation to one of understanding the structure of the space of examples.

Typeset by AMS-TEX